# Collaborative Unlabeled Data Optimization

### v1

@@ Conclusion.
In this paper, we introduce CoOpt, a pioneering data-centric, efficient, and parallelized framework for collaborative optimization of unlabeled data. This data-centric approach results in architecture-agnostic optimized data that are reusable across diverse network architectures, while simultaneously reducing the number of training iterations required, thereby enhancing overall computational efficiency. The collaborative mechanism effectively exploits distributed computation to substantially lower the cost per participant. Furthemore, within CoOpt, we identify a critical issue: Target Distribution Inconsistency, which arises from the diversity of prior models used in data optimization. To mitigate this, we propose a target alignment strategy. Extensive experiments conducted across various real-world scenarios demonstrate the superior effectiveness and efficiency of the CoOpt framework across diverse datasets and architectures. One limitation is that the experiments are primarily conducted on image datasets, without extending to other domain data.
Future Work.
In the future, our objectives include the following: (1) Develop advanced methodologies to ensure strong privacy guarantees; (2) Conduct an in-depth exploration of continuous data optimization through iterative processes, enabling dynamic collaboration among participants. (3) Integrate diverse domain datasets into the optimization framework. (3) Explore more practical scenarios, such as participants optimizing their local data and then uploading it to the shared platform. This paradigm can broaden the platform’s functionalities, where organizations and individuals can contribute to and benefit from a globally data ecosystem.
@@ Appendix A Related Work
Self-supervised learning [7] aims to exploit the intrinsic relationships within unlabeled data. For example, InstDisc [34] uses instance discrimination as a pretext task. Building on this, CMC [30] proposes to use multiple views of an image as positive samples and take another one as the negative. MoCo [15] significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples. SimCLR [6] highlights the importance of hard positive sample strategies by introducing data augmentation. Notably, BYOL [13] discards negative sampling and surpasses the performance of SimCLR [6].
@@
Knowledge distillation [17] employs soft labels generated by teacher models to improve the performance of a student model and expedite its training [11]. Many following works aim to enhance the use of soft labels for more effective knowledge transfer. For example, WSLD [45] analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off. DKD [44] decouples the logits and assigns different weights for the target and non-target classes. Moreover, several studies [37, 11] demonstrate that knowledge distillation can accelerate the training process.
@@
Dataset distillation [33] aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost. The majority of methods focus on optimizing images, which can be categorized into five primary approaches [22]: meta-learning frameworks [33, 46], gradient matching [43, 41], distribution matching [42, 38], trajectory matching [5, 14], and decoupling frameworks [38, 29]. Recently, some studies [27, 25] have shifted focus towards label distillation, aiming to obtain high-quality soft labels. This approach has demonstrated notable efficiency and effectiveness.
Summary. Both approaches are computationally expensive as they require task-specific models.
@@ B.1 Implementation Details
We utilize 10 pre-trained prior models, which are downloaded using the torchvision package. These models consist of 5 ImageNet-1K-based models, varying from small-scale to large-scale architectures, and 5 CLIP-based model. Unlabeled data is equally split among participants.
B.3 Diverse Network for Model training in Optimized data.
Table 5: Comparison of CoOpt with BYOL for Diverse Network on CIFAR-100.
Method
Network BYOL CoOpt
ResNet-50 60.4 $\pm$ 0.1 63.8 $\pm$ 0.0
ResNet-101 61.5 $\pm$ 0.2 65.7 $\pm$ 0.2
MobileNet-v2 24.0 $\pm$ 0.5 58.1 $\pm$ 0.0
Efficientnet-b0 2.3 $\pm$ 0.2 70.7 $\pm$ 0.2
ViT 38.5 $\pm$ 0.1 57.8 $\pm$ 0.1
We conduct experiments across various network architectures, comparing our approach, CoOpt, with BYOL, as presented in Table 5. Notably, CoOpt consistently demonstrates significant performance improvements over BYOL across multiple architectures. Furthermore, BYOL exhibits markedly inferior performance when applied to lightweight architectures such as MobileNet-v2 and EfficientNet-b0, in contrast to its performance with larger backbones like ResNet-50. One plausible reason for this performance drop is the instability of batch normalization (BN) statistics in the early layers of these lightweight networks. Specifically, the narrow layers (i.e., those with small channel width) in early stages produce activations with limited sample size per channel, resulting in high variance in the estimated batch mean and variance. Since BYOL relies on two differently augmented views of the same image, this BN-induced instability leads to exaggerated discrepancies between the two projected features, which deviates from the intended invariance objective of the method. This mismatch potentially impairs the optimization process and harms representation quality.
B.4 Diverse Prior Datasets
@@
B.5 Continuous Data Optimization
@@
The overall process is illustrated in Figure 6. The framework operates as a loop:

### v2

@@ 5 Conclusion
We introduce CoOpt, a pioneering data-centric, parallelized, and efficient framework for collaborative optimization of unlabeled data. This data-centric approach results in architecture-agnostic optimized data that are reusable across diverse network architectures, while simultaneously reducing the number of training iterations required, thereby enhancing overall efficiency. Furthemore, within CoOpt, we identify a critical issue: Target Distribution Inconsistency, which arises from the diversity of prior models used in data optimization. To mitigate this, we propose a lightweight target alignment strategy. Extensive experiments demonstrate the superior effectiveness and efficiency of the CoOpt framework across diverse datasets and architectures. One limitation is that when all prior models are extremely weak, the overall performance inevitably degrades. As future work, we aim to develop advanced strategies to more effectively exploit optimized data derived from all extremely weak priors, with our current study verifying efficiency in the early training stage.
@@ Appendix A Related Work
Self-supervised learning (Chen et al., 2020b) aims to exploit the intrinsic relationships within unlabeled data. For example, InstDisc (Wu et al., 2018) uses instance discrimination as a pretext task. Building on this, CMC (Tian et al., 2020) proposes to use multiple views of an image as positive samples and take another one as the negative. MoCo (He et al., 2020) significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples. SimCLR (Chen et al., 2020a) highlights the importance of hard positive sample strategies by introducing data augmentation. Notably, BYOL (Grill et al., 2020) discards negative sampling and surpasses the performance of SimCLR (Chen et al., 2020a).
@@
Knowledge distillation (Hinton, 2015) employs soft labels generated by teacher models to improve the performance of a student model and expedite its training (Dong et al., 2023). Many following works aim to enhance the use of soft labels for more effective knowledge transfer. For example, WSLD (Zhou et al., 2021) analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off. DKD (Zhao et al., 2022) decouples the logits and assigns different weights for the target and non-target classes. Moreover, several studies (Yim et al., 2017; Dong et al., 2023) demonstrate that knowledge distillation can accelerate the training process.
@@
Dataset distillation (Wang et al., 2018) aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost. The majority of methods focus on optimizing images, which can be categorized into five primary approaches (Lei & Tao, 2023): meta-learning frameworks (Wang et al., 2018; Zhou et al., 2022), gradient matching (Zhao et al., 2020; Zhao & Bilen, 2021), distribution matching (Zhao & Bilen, 2023; Yin et al., 2023), trajectory matching (Cazenavette et al., 2022; Guo et al., 2024), and decoupling frameworks (Yin et al., 2023; Sun et al., 2024). Recently, some studies (Shang et al., 2025; Qin et al., 2024) have shifted focus towards label distillation, aiming to obtain high-quality soft labels. This approach has demonstrated notable efficiency and effectiveness.
@@ B.1 Implementation Details
We utilize 10 CLIP-based pre-trained prior models varying from small-scale to large-scale architectures, which are downloaded using the torchvision package. Unlabeled data is equally split among participants.
B.3 Diverse Prior Datasets
@@
B.4 Continuous Data Optimization
@@
The overall process is illustrated in Fig. 7. The framework operates as a loop:
@@
Table 7: Performance Under Three Selection Scenarios.
Dataset BYOL No align Best Medium Worse
CIFAR-100 51.7 $\pm$ 0.1 44.7 $\pm$ 0.0 65.3 $\pm$ 0.0 63.1 $\pm$ 0.0 60.1 $\pm$ 0.1
B.5 Various Alignment strageties
As shown in Fig. 6(d), even when the best prior model is mis-selected (i.e., prior models are aligned to a medium-performing prior or even the worst prior, represented by the purple and red lines, respectively), the overall performance still surpasses the “no align” baseline. Furthermore, we present the final performance in Tab. 7. The results suggest that our method continues to outperform the baseline methods under three selection scenarios and demonstrates robustness to prior model mis-selection.
Table 8: Influence of Shared Data Size on Large-Scale ImageNet-1K.
Dataset BYOL 0.001 0.1
ImageNet-1K 61.9 $\pm$ 0.1 68.8 $\pm$ 0.1 68.7 $\pm$ 0.0
B.6 Size of Shared Data
The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Sec. 3.4. The shared dataset is public and randomly selected. This eliminates privacy concerns and does not require matching any specific data distribution, thus avoiding issues of distribution mismatch or privacy leakage. Moreover, we further evaluated our method on the large-scale ImageNet-1K. As shown in Tab. 8, only 0.001% of the data samples are sufficient to achieve performance comparable to using larger amounts of data. This minimal requirement imposes negligible additional overhead, ensuring scalability to very large datasets.